What is the probability of having a baby boy?
The sex of a child is determined by male sperm. There are only two sexes, so there is an equal (50-50) chance of having a boy or girl.
30 percent.
What is the probability of flipping a coin 4 times and it landing on heads each time?
p(heads)= 0.5
p(heads)^4= 0.0625
What is the probability that dinosaur bones are in your backyard?
It depends on where you live. If you live in Utica, Illinois, the probability is close to zero. If you live in Cheyenne, Wyoming, the odds are upped considerably.
I assume you mean a matching pair. If so, the answer is 13.
Considering the worst-case scenario, she may pick all six left-handed brown gloves and then all six left-handed black gloves. In which case, the next glove would definitely be a right-handed (black or brown) glove giving a matching pair.
Even in complete darkness, Cathy might be able to discern the difference between a right and left-handed glove (mittens not so much). So, as written, the answer is two. If we assume a matching pair, the answer is eight.
The young Couple to light up there house in the winter evenings.
The Poet To just look at its beauty to inspire him in writing poems.
If you have a 52 card deck and deal 48 cards, the probability that the Jack of Spades
is in the remaining 4 cards is the same as in the case of dealing 4 cards from the 52
card deck, the probability the Jack of Spades is in those 4 cards.
To solve the problem you can see there are 4 possible ways that the Jack of Spades
can come up in the 4 deal-led cards.
I.- It can come in the first card. P1= 1/52.
II.- Come in the second card. P2= 51/52 (1/51).
III.- Come in the third card. P3= 51/52 (50/51) (1/50).
IV.- Or in the fourth card. P4= 51/52 (50/51) (49/50) (1/49).
The probability looked for is P = P1+P2+P3+P4 = 1/13
Business application of probability?
Probability (and probability based models) are used in business scenarios to make decisions under uncertainty. A good example is maintaining safety stocks of inventory when there is a probability that the demand for product might be higher than the supply. Another application is in financial markets where the returns are not certain so people use probabilities to predict outcomes and hedge against uncertainties.
75%.
Explanation: if both parents are carriers than their Gametes would be Dd x Dd, resulting in DD (1/4), Dd (2/4), and dd(1/4), meaning that there is a 25% chance that they will have a child that has the disorder.
DD = A normal child that lacks the recessive gene, therefore not a carrier
Dd = A normal child that carries the recessive gene
dd = A child with the disorder + carries the gene, hence the child's future offspring will carry it as well.
Source: Self/Basic Punnet Square
When flipping a coin, there are 2 possible outcomes. When flipping 3 coins there are 8 possible outcomes (2^3=8).
As for the situation described, there is only one way for it to not be true, if all the coins land on the same side. So either all heads or all tails.
This leaves 8-2=6 possible outcomes resulting in the above situation.
Therefore the probability of the given situation is 6/8 or 3/4=75%
How is probability used in real life?
=Probability is used in many ways.==For example:==* gambling==*bettting odds==and anywhere in the world!=
Concept of Probability sampling and chi square test?
A probability sampling method is any method of sampling that utilizes some form of random selection. See: http://www.socialresearchmethods.net/kb/sampprob.php The simple random sample is an assumption when the chi-square distribution is used as the sampling distribution of the calculated variance (s^2). The second assumption is that the particular variable is normally distributed. It may not be in the sample, but it is assumed that the variable is normally distributed in the population. For a very good discussion of the chi-square test, see: http://en.wikipedia.org/wiki/Pearson%27s_chi-square_test
How can you relate probability and punnett squares?
The probability of an event is easily calculated by dividing the number of squares that predict the event by the total number of squares. For example, if a man and woman were having a child, and the man was tall with brown hair and the woman short with blonde hair, the probability of the child being, say, tall with blonde hair, would be 1(number of squares with tall/blonde as a result) divided by 4(total number of squares), which is .25, or 25%
What school runs the Great Penny toss?
The Ralph Bunche Computer Minischool in New York City ran the Great Penny toss.
What situations in which working with others can achieve positive outcomes?
give your ideas in a discussion or a meeting!! Hiza
The answer is dependent on whether of not you replace the marbles in the jar. If you do, the probability of drawing a red marble is 9 in 15 or 60%, every time.
If you do not replace the marbles, the probability of drawing a red marble is 2 in 8 or 25%.
What is classical probability?
Classical probability theory is concerned with carrying out probability calculations based on equally
likely outcomes. That is, it is assumed that the sample space has been constructed in such a way
that every subset of the sample space consisting of a single element has the same probability. If
the sample space contains n possible outcomes (#S = n), we must have for all s 2 S,
P(fsg) =
1
n
and hence for all E S
P(E) =
#E
n
:
More informally, we have
P(E) =
number of ways E can occur
total number of outcomes
:
An odd number is an integer of the form 2*n+1 where n is an integer.
Equivalently, an odd number is an integer which leaver a remainder of 1 when divided by 2.
It is 1/6*1/6 = 1/36.
If I were to flip a coin 4 times. What is the probability of me getting heads twice?
If the coin is not biased, the answer is 0.375